OpenStudy (anonymous):

Which property does the statement illustrate? Column A Column B 1. (n + 6) • 7 = 7 • (n + 6) 2. 3 • (4 • 9) = (3 • 4) • 9 3. (2x • 5y) • 8 = 2x • (5y • 8) 4. 5 • 3 = 3 • 5 A. commutative property of multiplication B. associative property of multiplication will fan if not already fanned

OpenStudy (anonymous):

i got u hang on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

this is the communitive property a × b = b × a

OpenStudy (anonymous):

this is the associative property (a × b) × c = a × (b × c)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now what do u think the answer is

OpenStudy (anonymous):

which equation matches the question?

OpenStudy (anonymous):

ohhhhh i got it now thanks for the help

OpenStudy (anonymous):

np